Given that U=f(q1, q2) and p1=2$, p2= 5$ per kg. If the budget constraint of a consumer is 100$, find the maximum quantity of goods q1 and q2 which the consumer can consume to maximise utility


1
Expert's answer
2021-03-14T19:53:14-0400

The budget constraint of the consumer can be written as follows:


"B=p_1q_1+p_2q_2,""100=2q_1+5q_2."

Let's find the Marginal Rate of Substitution (MRS):


"MRS=\dfrac{p_1}{p_2}=\dfrac{2}{5}=0.4"

Let's consider a special case of the "Cobb-Douglas" utility function, which has the form:


"U(q_1,q_2)=q_1^aq_2^b,"

where "a" and "b" are two constants. In this case the marginal rate of substitution (MRS) for the Cobb-Douglas utility function is:


"MRS=(\dfrac{a}{b})(\dfrac{q_2}{q_1})"

regardless of the values of "a" and "b".

Therefore, we can write:


"MRS=\dfrac{p_1}{p_2}=0.4=\dfrac{q_2}{q_1},""q_2=0.4q_1."

Therefore, the budget constraint becomes:


"100=2q_1+5\cdot0.4q_1=4q_1,""q_1=\dfrac{100}{4}=25."

Finally, we can find "q_2":


"q_2=0.4q_1=0.4\cdot25=10."

Answer:

"q_1=25, q_2=10."